OpenStudy (anonymous):

LCD of 1/t - 1/t^2+t ?

OpenStudy (zzr0ck3r):

\(\frac{1}{t}-\frac{1}{t^2+t}\)?

OpenStudy (imstuck):

\[\frac{ 1 }{ t }-\frac{ 1 }{ t ^{2} }+t\]Which is it? These are two very different things!

OpenStudy (anonymous):

the first one

OpenStudy (imstuck):

I recommend factoring a t out of the second denominator there and see if you notice anything that way.

OpenStudy (imstuck):

It's quite apparent when you do that.

OpenStudy (zzr0ck3r):

\(\frac{1}{t}-\frac{1}{t^2+t}=\frac{1}{t}-\frac{1}{t(t+1)}\) one term had t, the other haas t(t+1). so its lcm = t(t+1)

OpenStudy (anonymous):

yah that's what I did... btw thanks :)